How many liters each of a 30% antifreeze solution and a 60% antifreeze solution must be mixed to make 10 L of a 36% anti-freeze solution?
Answer by rothauserc(4718)

We have two equations to work with, let x be number of liters(L) of .30 solution, y be the number of L of .60 solution and 10*.36 = 3.6, then
x + y = 10
.30x + .60y = 3.60
solve first equation for x
x = 10 - y
substitute for x in second equation
.30(10-y) + .60y = 3.60
3-.30y +.60y = 3.60
.30y = .60
y = 2
therefore we have 2L of the .60 solution and 8L of the .30 solution